A 0.015 kg marble sliding to the right at .225 m/s on a frictionless surface makes an elastic head on collision with a 0.0025 kg marble moving to the left at .18 m/s. After the collision, the first marble moves to the left at .18 m/s. find the velocity of the canoe after the collision.
Answer #1

using conservation of momentum

m1 u1 + m2 u2 = m1 v1 + m2 v2

15* 0.225 + 2.5* (- 0.18) = 15*( - 0.18) + 2.5* v2

v2 = 2.25 m/s

========

Comment before rate in case any doubt, will reply for sure.. goodluck